# On the well-posedness of SPDEs with locally Lipschitz coefficients

### v2

@@ Abstract.
where $u=u(t\,,x)$ is defined for $(t\,,x)\in(0\,,\infty)\times\mathbb{R}$, and $\dot{W}$ denotes space-time white noise. We prove that this SPDE is well posed solely under the assumptions that the initial condition $u(0)$ is bounded and measurable, and $b$ and $\sigma$ are locally Lipschitz continuous functions and have at most linear growth. Our method is based on a truncation argument together with moment bounds and tail estimates of the truncated solution. The results naturally generalize to the case where $b$ and $\sigma$ are time dependent with uniform-in-time growth and oscillation properties. Additionally, our method can be extended to the stochastic wave equation.
Keywords: SPDEs, space-time white noise, existence and uniqueness.
AMS 2010 subject classification: 60H15; 60H07, 60F05.
The research of the second author is supported in part by the United States’ National Science Foundation grants DMS-1855439 and DMS-2245242
The third author acknowledges support from the Spanish MINECO grant PID2022-138268NB-100 and Ayudas Fundacion BBVA a Proyectos de Investigación Científica 2021
@@ 1. Introduction
where $(t\,,x)\in(0\,,\infty)\times\mathbb{R}$, subject to $u(0\,,x)=u_{0}(x)$, and the forcing $\dot{W}$ is space-time white noise; that is, $\dot{W}$ is a generalized Gaussian random field with mean zero and $\operatorname{\mathrm{Cov}}[\dot{W}(t\,,x)\,,\dot{W}(s\,,y)]=\delta_{0}(t-s)\delta_{0}(x-y)$ for all $t,s\geq 0$ and $x,y\in\mathbb{R}$. It is well known that (1.1) is well posed when $b$ and $\sigma$ are Lipschitz in their spatial variable uniformly in their time variable; see for example Dalang [Dalang] and Walsh [Walsh]. We aim to extend this result to the setting where “Lipschitz” is replaced by “local Lipschitz with linear growth.”
@@
The present undertaking yields an infinite-dimensional version of one of the foundational results of the theory of stochastic differential equations (SDEs) which asserts that if $Y=\{Y_{t}\}_{t\geq 0}$ denotes a standard Brownian motion on $\mathbb{R}$ and $x_{0}\in\mathbb{R}$ is non random, then the one-dimensional Itô-type SDE $\mathrm{d}X_{t}=b(t\,,X_{t})\,\mathrm{d}t+\sigma(t\,,X_{t})\,\mathrm{d}Y_{t}$ ($t>0$) subject to $X_{0}=x_{0}$ has a unique solution provided only that $b(t\,,x)$ and $\sigma(t\,,x)$ are locally Lipschitz in $x$ with at-most linear growth, all valid uniformly in $t$. The preceding follows from the more classical result about SDEs with Lipschitz coefficients and a stopping time argument. See for example Exercise (2.10) of Revuz and Yor [RevuzYor, p. 383]. It is possible to use essentially the same method in order to extend the preceding to the multidimensional setting where $Y$ denotes a $d$-dimensional Brownian motion, $b:(0\,,\infty)\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\sigma:(0\,,\infty)\times\mathbb{R}^{d}\to(\mathbb{R}^{d})^{2}$.
@@
An infinite-dimensional extension appears in a forthcoming monograph by Robert C. Dalang and Marta Sanz-Solé [bookDS]. There, Dalang and Sanz-Solé show the well-posedness of (1.1), where instead $(t\,,x)\in(0\,,\infty)\times I$ in the case that $I\subset\mathbb{R}$ is a bounded interval (together with any of the usual boundary conditions). A key step in their analysis is the observation that if $b$ and $\sigma$ are locally Lipschitz in their space variable, uniformly in the time variable, then (1.3) below has a predictable random-field solution up to the stopping time $\tau(I)=\inf\{t>0:\sup_{x\in I}|u(t\,,x)|=\infty\}$ $[\inf\varnothing=\infty]$, and $\mathrm{P}\{\tau(I)>0\}=1$ when $I\subset\mathbb{R}$ is bounded. Such stopping-time arguments are typically not used to study equations on $\mathbb{R}_{+}\times\mathbb{R}$ because one expects the resulting solution, if there is one, to be unbounded with probability one. Therefore, in order to produce solutions to the SPDE (1.1) one needs to introduce a different argument.
@@
The purpose of this article is to provide one such argument: We use truncation, as has been done previously, but replace the stopping-time argument by pointwise tail probability estimates for the truncated solution. In this way we are able to show that, with high probability, the truncated solution is not large, uniformly in the truncation level. At that stage, our method boils down to an application of an amusing elementary lemma on summability [Lemma 7] that plays an essential role in this work.
It is also worth noting that the techniques developed in this paper extend beyond the stochastic heat equation (1.1). Specifically, similar methods can be applied to the stochastic wave equation of the following form,
$\partial^{2}_{t}u(t\,,x)=\tfrac{1}{2}\partial^{2}_{x}u(t\,,x)+b(t\,,u(t\,,x))+\sigma(t\,,u(t\,,x))\dot{W}(t\,,x),$
where $t\geq 0$, $x\in\mathbb{R}$, subject to appropriate initial conditions, see [bookDS, Chapter 4] for the existence and uniqueness under globally Lipschitz condition. Although the stochastic wave equation presents additional challenges due to its hyperbolic nature, the local Lipschitz and linear growth assumptions on $b$ and $\sigma$ allow our approach to handle the existence and uniqueness of solutions in this setting as well. The key adjustments consist of utilizing the wave kernel in place of the heat kernel and establishing similar tail probability estimates for the truncated solutions.
@@
Assumption 2.
The functions $b:(0\,,\infty)\times\mathbb{R}\rightarrow\mathbb{R}$ and $\sigma:(0\,,\infty)\times\mathbb{R}\rightarrow\mathbb{R}$ are locally Lipschitz continuous in their space variable with at most linear growth, uniformly in their time variable. In other words, $0<\text{\rm Lip}_{n}(b),\text{\rm Lip}_{n}(\sigma)<\infty$ and $0\leq L_{b},L_{\sigma}<\infty,$ for all real numbers $n>0$ where, for every space-time function $\psi$,
@@
where, for every space-time function $\psi$,
$\mathcal{I}_{\psi}(t\,,x)=\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)\psi(s\,,u(s\,,y))\,\mathrm{d}s\,\mathrm{d}y,$ (1.4)
the symbol “*” denotes convolution, and
$p_{r}(z)=\frac{\exp\{-z^{2}/(2r)\}}{\sqrt{2\pi r}}\qquad\text{for all $r>0$ and $z\in\mathbb{R}$}.$
The following is the main result of this paper.
Theorem 3.
Under Assumptions 1 and 2, (1.1) has a unique random-field solution that satisfies the following:
$\adjustlimits{\sup}_{t\in[0,T]}{\sup}_{x\in\mathbb{R}}\mathrm{E}\left(|u(t\,,x)|^{k}\right)<\infty\quad\text{for all $T>0$ and $k\geq 1$}.$
Let us end the Introduction with a brief description outline of the paper. In §2 we introduce the truncated solution and develop some moment and tail estimates. The remaining details of the proof of Theorem 3 are gathered in §3, and use the earlier results of the paper.
Throughout this paper, we write $\|X\|_{p}=\{\mathrm{E}(|X|^{p})\}^{1/p}$ for all $p\geq 1$ and $X\in L^{p}(\Omega)$. For every space-time function $f:(0\,,\infty)\times\mathbb{R}\to\mathbb{R}$, $\text{\rm Lip}(f)$ denotes the optimal Lipschitz constant of $f$; that is,
$\text{\rm Lip}(f)={\color[rgb]{0,0,0}\sup_{t>0}\sup_{\begin{subarray}{c}a,b\in\mathbb{R}\\ a\neq b\end{subarray}}\frac{|f(t\,,b)-f(t\,,a)|}{|b-a|}.}$ (1.5)
In particular, $f$ is globally Lipschitz continuous in $x$, uniformly in $t$, iff $\text{\rm Lip}(f)<\infty$. If $f$ depends only on a spatial variable $x$, then $\text{\rm Lip}(f)$ still makes sense provided that we extend $f$ to a space-time function as follows $f(t\,,x)=f(x)$, in the usual way.
@@ 2. Truncation, and preliminary results
In this section we describe the truncation argument that was referred to verbally in the Introduction. Define,
and, for every real number $N>0$, two functions $b_{N}:(0\,,\infty)\times\mathbb{R}\to\mathbb{R}$ and $\sigma_{N}:(0\,,\infty)\times\mathbb{R}\to\mathbb{R}$ as follows: For all $t>0$ and $\psi:\mathbb{R}_{+}\times\mathbb{R}\to\mathbb{R}$ let
Choose and fix an arbitrary $N>0$ and note that $b_{N}$ and $\sigma_{N}$ are globally Lipschitz. In fact, recall (1.5) in order to see that $\text{\rm Lip}(b_{N})=L_{N,b}$ and $\text{\rm Lip}(\sigma_{N})=L_{N,\sigma}$, and
@@
uniformly in $N>0$. Thanks to standard theory [Dalang, Walsh], the following SPDE has a predictable mild solution: For $(t\,,x)\in(0\,,\infty)\times\mathbb{R}$,
@@
subject to $u_{N}(0\,,x)=u_{0}(x)$. Moreover this solution is unique subject to the condition,
$\adjustlimits{\sup}_{t\in(0,T)}{\sup}_{x\in\mathbb{R}}\mathrm{E}\left(|u_{N}(t\,,x)|^{k}\right)<\infty\quad\text{for all $N,T>0$ and $k\geq 1$}.$ (2.3)
See Dalang and Sanz-Solé [bookDS, Theorem 4.2.1]. As usual, (2.2) is short-hand for the random integral equation,
$u_{N}(t\,,x)=(p_{t}*u_{0})(x)+\mathcal{I}^{N}_{b}(t\,,x)+\mathcal{I}^{N}_{\sigma}(t\,,x),$ (2.4)
where $\mathcal{I}^{N}_{b}$ and $\mathcal{I}^{N}_{\sigma}$ are the truncated analogous of the integrals in (1.4), where for every space-time function $\psi$,
$\mathcal{I}^{N}_{\psi}(t\,,x)=\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)\psi_{N}(s\,,u_{N}(s\,,y))\,\mathrm{d}s\,\mathrm{d}y.$ (2.5)
@@
$\adjustlimits{\sup}_{N>0}{\sup}_{x\in\mathbb{R}}\mathrm{E}\left(|u_{N}(t\,,x)|^{k}\right)\leq 4^{k}(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+1)^{k}\mathrm{e}^{128L_{\sigma}^{4}k^{3}t}.$
@@
where $I_{1}=\|\mathcal{I}^{N}_{b}(t\,,x)\|_{k}$ and $I_{2}=\|\mathcal{I}^{N}_{\sigma}(t\,,x)\|_{k}.$ We begin by estimating $I_{1}$ as follows. Thanks to (1.2) and the Minkowski inequality and the fact that $\|b_{N}(s\,,u_{N}(s\,,y))\|_{k}\leq L_{b}(1+\|u_{N}(s\,,y)\|_{k})$,
@@
$I_{1}\leq L_{b}\left[t+\mathcal{N}_{k,\beta}(u_{N})\int_{0}^{t}\mathrm{e}^{\beta s}\,\mathrm{d}s\right]\leq L_{b}\left[t+\frac{\mathrm{e}^{\beta t}}{\beta}\mathcal{N}_{k,\beta}(u_{N})\right].$
@@
$I_{1}\leq\frac{L_{b}\mathrm{e}^{\beta t}}{\beta}\left[1+\mathcal{N}_{k,\beta}(u_{N})\right].$ (2.8)
@@
Thanks to (1.2) and the fact that $\|\sigma_{N}(s\,,u_{N}(s\,,y))\|\leq 2L_{\sigma}^{2}(1+\|u_{N}(s\,,y)\|_{k}^{2})$,
@@
$\|p_{r}\|_{L^{2}(\mathbb{R})}^{2}=(p_{r}*p_{r})(0)=p_{2r}(0)=\frac{1}{2\sqrt{\pi r}}\qquad\text{for every $r>0$}.$ (2.10)
@@
$I_{2}^{2}\leq\frac{4k\mathrm{e}^{2\beta t}L_{\sigma}^{2}}{\sqrt{2\beta}}\left(1+\left[\mathcal{N}_{k,\beta}(u_{N})\right]^{2}\right).$
@@
$I_{2}\leq\frac{2\sqrt{k}\mathrm{e}^{\beta t}L_{\sigma}}{(2\beta)^{1/4}}\left(1+\mathcal{N}_{k,\beta}(u_{N})\right).$ (2.11)
@@
$\|u_{N}(t\,,x)\|_{k}$ $\leq\|u_{0}\|_{L^{\infty}(\mathbb{R})}+\frac{L_{b}\mathrm{e}^{\beta t}}{\beta}\left[1+\mathcal{N}_{k,\beta}(u_{N})\right]+\frac{2\sqrt{k}\mathrm{e}^{\beta t}L_{\sigma}}{(2\beta)^{1/4}}\left(1+\mathcal{N}_{k,\beta}(u_{N})\right)$
$\leq\|u_{0}\|_{L^{\infty}(\mathbb{R})}+\mathrm{e}^{\beta t}\left(\frac{L_{b}}{\beta}+\frac{2\sqrt{k}L_{\sigma}}{(2\beta)^{1/4}}\right)\left(1+\mathcal{N}_{k,\beta}(u_{N})\right).$
@@
$\mathcal{N}_{k,\beta}(u_{N})\leq\|u_{0}\|_{L^{\infty}(\mathbb{R})}+\left(\frac{L_{b}}{\beta}+\frac{2\sqrt{k}L_{\sigma}}{(2\beta)^{1/4}}\right)\left(1+\mathcal{N}_{k,\beta}(u_{N})\right).$
@@
Proposition 4 omits the case that $\sigma$ is a constant. The following takes care of that case, and at the same time improves on Proposition 4 when $\sigma$ is bounded but $\text{Lip}(\sigma)>0$.
Proposition 5.
If $\sigma:(0\,,\infty)\times\mathbb{R}\to\mathbb{R}$ is bounded, then
@@
We modify the proof of Proposition 4 by first observing that (2.6) remains valid, and so does (2.8). We start with (2.9) and estimate $I_{2}$, using (2.10), simply as follows:
$I_{2}^{2}$ $\leq 4k\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}^{2}\int_{0}^{t}\|p_{r}\|_{L^{2}(\mathbb{R})}^{2}\,\mathrm{d}r=\frac{2k}{\sqrt{\pi}}\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}^{2}\int_{0}^{t}\frac{\mathrm{d}r}{\sqrt{r}}$
@@
$\|u_{N}(t\,,x)\|_{k}\leq\|u_{0}\|_{L^{\infty}(\mathbb{R})}+\frac{L_{b}\exp(\beta t)}{\beta}\left[1+\mathcal{N}_{k,\beta}(u_{N})\right]+2\sqrt{k}\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}t^{1/4}.$
Since the right-hand side does not depend on $(t\,,x)$, we divide by $\exp(\beta t)$ and optimize over $(t\,,x)$ to find that
$\mathcal{N}_{k,\beta}(u_{N})\leq\|u_{0}\|_{L^{\infty}(\mathbb{R})}+2\sqrt{k}\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}t^{1/4}+\frac{L_{b}}{\beta}\left[1+\mathcal{N}_{k,\beta}(u_{N})\right],$
@@
$\mathrm{P}\left\{|u_{N+1}(t\,,x)|\geq\mathrm{e}^{N}\right\}\leq\exp\left(-\frac{N^{3/2}}{(512t)^{1/2}L_{\sigma}^{2}}\right),$

### v3

@@ Abstract.
[…]
Key words and phrases:
SPDEs, space-time white noise, existence and uniqueness
2010 Mathematics Subject Classification:
60H15; 60H07, 60F05
Research supported by the Leverhulme Trust Fellowship IF-2025-040, the US-NSF grants DMS-1855439 and DMS-2245242, the Spanish MINECO grant PID2022-138268NB-100, and Ayudas Fundacion BBVA a Proyectos de Investigación Científica 2021
@@ 1. Introduction
[…]
$\operatorname{\mathrm{Cov}}[\dot{W}(t\,,x)\,,\dot{W}(s\,,y)]=\delta_{0}(t-s)\delta_{0}(x-y)\quad\text{for all }t,s\geq 0\text{ and }x,y\in\mathbb{R}.$
@@
It is well known that (1.1) is well posed when $b$ and $\sigma$ are Lipschitz in their spatial variable uniformly in their time variable; see for example Dalang [Dalang] and Walsh [Walsh]. We aim to extend this result to the setting where “Lipschitz” is replaced by “local Lipschitz with linear growth.”
The present undertaking yields an infinite-dimensional version of one of the foundational results of the theory of stochastic differential equations (SDEs) which asserts that if $Y=\{Y_{t}\}_{t\geq 0}$ denotes a standard Brownian motion on $\mathbb{R}$ and $x_{0}\in\mathbb{R}$ is non random, then the one-dimensional Itô-type SDE
$\mathrm{d}X_{t}=b(t\,,X_{t})\,\mathrm{d}t+\sigma(t\,,X_{t})\,\mathrm{d}Y_{t}\qquad(t>0)$
@@
subject to $X_{0}=x_{0}$ has a unique solution provided only that b(t,x)b(t\,,x) and $\sigma(t\,,x)$ are locally Lipschitz in $x$ with at-most linear growth, all valid uniformly in $t$. The preceding follows from the more classical result about SDEs with Lipschitz coefficients and a stopping time argument. See for example Exercise (2.10) of Revuz and Yor [RevuzYor, p. 383]. It is possible to use essentially the same method in order to extend the preceding to the multidimensional setting where $Y$ denotes a $d$-dimensional Brownian motion, $b:(0\,,\infty)\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ and $\sigma:(0\,,\infty)\times\mathbb{R}^{d}\to(\mathbb{R}^{d})^{2}$.
@@
An infinite-dimensional extension appears in a forthcoming monograph by Robert C. Dalang and Marta Sanz-Solé [bookDS]. There, Dalang and Sanz-Solé show the well-posedness of (1.1), where instead $(t\,,x)\in(0\,,\infty)\times I$ in the case that $I\subset\mathbb{R}$ is a bounded interval (together with any of the usual boundary conditions). A key step in their analysis is the observation that if $b$ and $\sigma$ are locally Lipschitz in their space variable, uniformly in the time variable, then (1.3) below has a predictable random-field solution up to the stopping time
$\tau(I)=\inf\{t>0:\sup_{x\in I}|u(t\,,x)|=\infty\}\qquad[\inf\varnothing=\infty],$
and $\mathrm{P}\{\tau(I)>0\}=1$ when $I\subset\mathbb{R}$ is bounded. Such stopping-time arguments are typically not used to study equations on $\mathbb{R}_{+}\times\mathbb{R}$ because one expects the resulting solution, if there is one, to be unbounded with probability one. Therefore, in order to produce solutions to the SPDE (1.1) one needs to introduce a different argument.
@@
The purpose of this article is to provide one such argument: We use truncation, as has been done previously, but replace the stopping-time argument by pointwise tail probability estimates for the truncated solution. In this way we are able to show that, for a wide family of locally Lipschitz functions $b$ and $\sigma$ of linear growth, and with high probability, the truncated solution is not large, uniformly in the truncation level.
Our method is elementary as it uses only standard SPDE estimates, and more significantly works without need for any unnecessary technical assumptions. But we are quick to mention that other potential approaches to such problems already exist in the literature: It might be for example possible to adapt methods of finite-dimensional SDEs, such as weak existence (in the probabilistic sense) followed by strong uniqueness via a Yamada-Watanabe (see Kurtz [K14]) or Gyöngy-Krylov (see Gyöngy and Krylov [GK96]) type arguments. In the infinite-dimensional context of SPDEs, weak existence for (1.1) was established in Shiga [S94, Theorem 2.6] when additionally $b(0)\geq 0$ and $\sigma(0)=0$. And Mytnik et al [MPS06, Theorem 1.2] establish the weak existence for (1.1) when $b\equiv 0$, $\sigma=$ Hölder continuous with linear growth, all forced by high-dimensional, spatially correlated noise. There is also a method that involves transforming the SPDE into a random integral equation which can in turn be compared with a deterministic PDE; see Salins [Salins]. This method does not lend itself easily to the case that $\sigma$ is non constant unless there are additional contraints such as $b(0)=0$, $\sigma(0)=0$, and more stringent conditions on the initial function; see Chen et al [CFHS] for the current state of the art of that method.
@@
Assumption 1.2.
[…]
@@
where the symbol “*” denotes convolution,
$p_{r}(z)=(2\pi r)^{-1/2}\exp\{-z^{2}/(2r)\}\qquad\text{for all $r>0$ and $z\in\mathbb{R}$},$
and $\mathcal{I}_{b}$ and $\mathcal{I}_{\sigma}$ are the following random fields,
$\begin{split}\mathcal{I}_{b}(t\,,x)&=\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)b(s\,,u(s\,,y))\,\mathrm{d}s\,\mathrm{d}y,\\ \mathcal{I}_{\sigma}(t\,,x)&=\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)\sigma(s\,,u(s\,,y))\,W(\mathrm{d}s\,\mathrm{d}y),\end{split}$ (1.4)
where the second (stochastic) integral is understood in the sense of Walsh [Walsh].
[…]
We will need the following assumption on the Lipschitz coefficients of $b_{N}$ and $\sigma_{N}$. Set
Assumption 1.3.
If $L_{\sigma}>0$ then we assume that
$L_{N,\sigma}=\mathscr{o}(N^{3/8})\quad\text{and}\quad L_{N,b}/L_{N,\sigma}^{4}=\mathcal{O}(1)\qquad\text{as $N\to\infty$}.$ (1.6)
If $\sigma$ is bounded, then we assume that
$L_{N,\sigma}=\mathscr{o}(\mathrm{e}^{N/2})\quad\text{and}\quad L_{N,b}/L_{N,\sigma}^{4}=\mathcal{O}(1)\qquad\text{as $N\to\infty$}.$ (1.7)
It has been widely believed for a long time that Theorem 1.4 ought to hold solely under the assumptions of locally Lipschitz coefficients $b$ and $\sigma$ with linear growth. The following is the main result of this paper, and makes some partial success toward the resolution of this long-standing problem.
Theorem 1.4.
If $(b\,,\sigma)$ satisfy Assumption 1.3, then (1.1) has a unique random-field solution that satisfies the following:
$\sup_{t\in[0,T]}\sup_{x\in\mathbb{R}}\mathrm{E}\left(|u(t\,,x)|^{k}\right)<\infty\quad\text{for all $T>0$ and $k\geq 1$}.$
Next we make some remarks in order to highlight the limitations and scope of Theorem 1.4 and its proof.
Remark 1.5.
The condition $L_{N,b}/L_{N,\sigma}^{4}=\mathcal{O}(1)$ has new content only when $L_{N,\sigma}\to\infty$ as $N\to\infty$. Indeed, if $\sup_{N}L_{N,\sigma}<\infty$, then the first stated condition ensures that both $b$ and $\sigma$ are globally Lipschitz continuous.
Remark 1.6.
Condition (1.6) is quite restrictive as it necessarily implies that $\text{\rm Lip}_{n}(\sigma)=\mathscr{o}(|\log n|^{3/8})$ and $\text{\rm Lip}_{n}(b)=\mathscr{o}(|\log n|^{3/2})$ as $n\to\infty$; see (1.5).
Remark 1.7 (Oscillatory diffusion coefficients).
Theorem 1.4 has stronger content when $\sigma$ is bounded for then (1.7) allows for a wide class of drifts $b$ when $\sigma$ is highly oscillatory. First, note that the first condition in (1.7) is equivalent to the statement that $\text{\rm Lip}_{n}(\sigma)=\mathscr{o}(\sqrt{n})$ as $n\to\infty$. Now consider an oscillatory diffusion coefficient such as $\sigma(x)=\sin(1000(1+|x|)^{1/4})$ for all $x\in\mathbb{R}$; see Figure 1. Then, $\sigma$ satisfies (1.7) together with every drift function that satisfies $\text{\rm Lip}_{n}(b)=\mathscr{o}(n)$.
[…]
Figure 1. An example of how Assumption 1.3 is less restrictive for the drift when $\sigma$ fluctuates wildly
Remark 1.8.
The method of proof of Theorem 1.4 allows for minor improvements of the first parts of (1.6) and (1.7). For example, the first condition in (1.6) can be improved slightly to $L_{N,\sigma}=\mathscr{o}(N^{2/3})$ by, instead of choosing the parameter $k$ as in (3.9), choosing it as $k(N)=A_{1}(N/T)^{1/2}L_{N,\sigma}^{-4/3}$ for a suitably small constant $A_{1}$ that is independent of $N$ and $T$. Then one obtains the main part of the proof – that is (3.1) below – from the fact that $\|u_{N+1}(t\,,x)-u_{N}(t\,,x)\|_{k}\leq\|u_{N+1}(t\,,x)-u_{N}(t\,,x)\|_{k(N)}$ for all large $N$ [Jensen’s inequality]. We have omitted the details of such improvements as they appear to be primarily technical in nature.
Let us end the Introduction with a brief outline of the paper. In §2 we introduce the truncated solution and develop some moment and tail estimates. The remaining details of the proof of Theorem 1.4 are gathered in §3, and use the earlier results of the paper.
[…]
$\text{\rm Lip}(f)=\sup_{t>0}\sup_{a,b\in\mathbb{R}:a\neq b}|f(t\,,b)-f(t\,,a)|/|b-a|.$
In particular, $f$ is globally Lipschitz continuous in $x$, uniformly in $t$, iff $\text{\rm Lip}(f)<\infty$. If $f$ depends only on a spatial variable $x$, then $\text{\rm Lip}(f)$ still makes sense provided that we extend $f$ to a space-time function as follows f(t,x)=f(x)f(t\,,x)=f(x), in the usual way.
@@ 2. Truncation, and preliminary results
In this section, we give more information about the truncation argument referred to verbally in the introduction. We begin by recalling the definition of $b_{N}$ and $\sigma_{N}$ and note that they are globally Lipschitz functions. In fact,
@@
[…]
@@
subject to $u_{N}(0\,,x)=u_{0}(x)$. Moreover this solution is unique subject to
$\sup_{t\in(0,T)}\sup_{x\in\mathbb{R}}\mathrm{E}\left(|u_{N}(t\,,x)|^{k}\right)<\infty\quad\text{for all $N,T>0$ and $k\geq 1$}$
and jointly continuous in $L^{k}(\Omega)$. See Dalang and Sanz-Solé [bookDS, Theorems 4.2.1 and 4.2.8]. As usual, (2.1) is short-hand for the random integral equation,
$u_{N}(t\,,x)=(p_{t}*u_{0})(x)+\mathcal{I}^{N}_{b_{N}}(t\,,x)+\mathcal{I}^{N}_{\sigma_{N}}(t\,,x),$ (2.2)
where $\mathcal{I}^{N}_{b_{N}}$ and $\mathcal{I}^{N}_{\sigma_{N}}$ are the truncated analogous of the integrals in (1.4) and are given by
$\begin{split}\mathcal{I}^{N}_{b_{N}}(t\,,x)&=\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)b_{N}(s\,,u_{N}(s\,,y))\,\mathrm{d}s\,\mathrm{d}y,\\ \mathcal{I}_{\sigma_{N}}^{N}(t\,,x)&=\int_{(0,t)\times\mathbb{R}}p_{t-s}(y-x)\sigma_{N}(s\,,u_{N}(s\,,y))\,W(\mathrm{d}s\,\mathrm{d}y).\end{split}$ (2.3)
@@
$\sup_{N>0}\sup_{x\in\mathbb{R}}\mathrm{E}\left(|u_{N}(t\,,x)|^{k}\right)\leq 4^{k}(\|u_{0}\|_{L^{\infty}(\mathbb{R})}+1)^{k}\mathrm{e}^{128L_{\sigma}^{4}k^{3}t}.$
@@
where $I_{1}=\|\mathcal{I}^{N}_{b_{N}}(t\,,x)\|_{k}$ and $I_{2}=\|\mathcal{I}^{N}_{\sigma_{N}}(t\,,x)\|_{k}.$ We begin by estimating $I_{1}$ as follows. Thanks to (1.2) and the Minkowski inequality and the fact that
$\|b_{N}(s\,,u_{N}(s\,,y))\|_{k}\leq L_{b}(1+\|u_{N}(s\,,y)\|_{k}),$
we can see that
@@
$I_{1}\leq L_{b}\left[t+\mathcal{N}_{k,\beta}(u_{N})\int_{0}^{t}\mathrm{e}^{\beta s}\,\mathrm{d}s\right]\leq L_{b}\left[t+\beta^{-1}\mathrm{e}^{\beta t}\mathcal{N}_{k,\beta}(u_{N})\right].$
@@
$I_{1}\leq L_{b}\beta^{-1}\mathrm{e}^{\beta t}\left[1+\mathcal{N}_{k,\beta}(u_{N})\right].$ (2.6)
@@
Thanks to (1.2) and the fact that $\|\sigma_{N}(s\,,u_{N}(s\,,y))\|_{k}^{2}\leq 2L_{\sigma}^{2}(1+\|u_{N}(s\,,y)\|_{k}^{2})$,
@@
$\|p_{r}\|_{L^{2}(\mathbb{R})}^{2}=(p_{r}*p_{r})(0)=p_{2r}(0)=\tfrac{1}{2}(\pi r)^{-1/2}\qquad\text{for every $r>0$}.$ (2.8)
@@
$I_{2}^{2}\leq 4k\mathrm{e}^{2\beta t}L_{\sigma}^{2}\left(1+\mathcal{N}_{k,\beta}(u_{N})^{2}\right)/\sqrt{2\beta}.$
@@
$I_{2}\leq 2\sqrt{k}(2\beta)^{-1/4}\mathrm{e}^{\beta t}L_{\sigma}\left(1+\mathcal{N}_{k,\beta}(u_{N})\right).$ (2.9)
@@
$\|u_{N}(t\,,x)\|_{k}$
$\leq\|u_{0}\|_{L^{\infty}(\mathbb{R})}+L_{b}\beta^{-1}\mathrm{e}^{\beta t}\left[1+\mathcal{N}_{k,\beta}(u_{N})\right]+2\sqrt{k}(2\beta)^{-1/4}\mathrm{e}^{\beta t}L_{\sigma}\left(1+\mathcal{N}_{k,\beta}(u_{N})\right)$
$\leq\|u_{0}\|_{L^{\infty}(\mathbb{R})}+\mathrm{e}^{\beta t}\left(L_{b}\beta^{-1}+2\sqrt{k}L_{\sigma}(2\beta)^{-1/4}\right)\left(1+\mathcal{N}_{k,\beta}(u_{N})\right).$
@@
$\mathcal{N}_{k,\beta}(u_{N})\leq\|u_{0}\|_{L^{\infty}(\mathbb{R})}+\left(L_{b}\beta^{-1}+2\sqrt{k}L_{\sigma}(2\beta)^{-1/4}\right)\left(1+\mathcal{N}_{k,\beta}(u_{N})\right).$
@@
Remark 2.2.
The requirement that $k\geq\max(2,L_{b}^{1/2}L_{\sigma}^{-2})$ is a technical consequence of the proof. We can always choose, without incurring loss of generality, $L_{\sigma}$ large enough to ensure that $L_{b}^{1/2}L_{\sigma}^{-2}\leq 2$, in order to ensure that the result stated in the proposition holds for all $k\geq 2$.
The conclusion of Proposition 2.1 can be improved when $\sigma$ is constant. The following takes care of that case, and at the same time improves Proposition 2.1 when $\sigma$ is bounded but $\text{Lip}(\sigma)>0$.
Proposition 2.3.
[…]
@@
We modify the proof of Proposition 2.1 by first observing that (2.4) remains valid, and so does (2.6). We start with (2.7) and estimate $I_{2}$, using (2.8), simply as follows:
$I_{2}^{2}$ $\leq 4k\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}^{2}\int_{0}^{t}\|p_{r}\|_{L^{2}(\mathbb{R})}^{2}\,\mathrm{d}r=2k\pi^{-1/2}\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}^{2}\int_{0}^{t}\mathrm{d}r/\sqrt{r}$
@@
$\|u_{N}(t\,,x)\|_{k}\leq\|u_{0}\|_{L^{\infty}(\mathbb{R})}+L_{b}\beta^{-1}\exp(\beta t)\left[1+\mathcal{N}_{k,\beta}(u_{N})\right]+2\sqrt{k}\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}t^{1/4}.$
Since the right-hand side does not depend on (t,x)(t\,,x), we divide by $\exp(\beta t)$ and optimize over (t,x)(t\,,x) to find that
$\mathcal{N}_{k,\beta}(u_{N})\leq\|u_{0}\|_{L^{\infty}(\mathbb{R})}+2\sqrt{k}\|\sigma\|_{L^{\infty}(\mathbb{R}_{+}\times\mathbb{R})}t^{1/4}+L_{b}\beta^{-1}\left[1+\mathcal{N}_{k,\beta}(u_{N})\right],$
@@
$\mathrm{P}\left\{|u_{N+1}(t\,,x)|\geq\mathrm{e}^{N}\right\}\leq\exp\left(-\frac{N^{3/2}}{64L_{\sigma}^{2}\sqrt{t}}\right),$ (2.10)